OpenStudy (mysticfury):

does this function have a negative or positive rate of change? 15x + 5y = −30

OpenStudy (mysticfury):

@Photon336

OpenStudy (igreen):

Convert to slope-intercept form. Do you know how?

OpenStudy (mysticfury):

nope @iGreen i imagined thats what i had to do

OpenStudy (photon336):

well for starters, think of it as any equation. you want to get y by itself how do you do that?

OpenStudy (mysticfury):

-5 from both sides? @Photon336

OpenStudy (photon336):

\[15x+5y = -30\]

OpenStudy (photon336):

We've got to get y by itself, we do this by subtracting 15 from both sides. \[5y = -30-15x \]

OpenStudy (photon336):

@MysticFury

OpenStudy (mysticfury):

oh okay so i was close

OpenStudy (photon336):

then what do we do?

OpenStudy (mysticfury):

5y = 15x?

OpenStudy (photon336):

so it's 5y = -30-15x

OpenStudy (photon336):

but we're not done yet

OpenStudy (mysticfury):

umm idk subtract?

OpenStudy (mysticfury):

@Photon336

OpenStudy (mysticfury):

@iGreen @UliNK @Michele_Laino

OpenStudy (mysticfury):

@KamiBug

OpenStudy (mysticfury):

please help @AloneS and @UliNK this is my last question

OpenStudy (ulink):

@MysticFury I honestly have no idea.

OpenStudy (mysticfury):

okay thanks anyway @UliNK

OpenStudy (ulink):

I wish I could be of some help, but I am not good at Mathematics.

OpenStudy (ulink):

Sorry ;(

OpenStudy (mysticfury):

i said its okay @UliNK

alones (alones):

Okay i don't really understand but i solved it out Let's solve for x. 15x+5y=−30 Step 1: Add -5y to both sides. 15x+5y+−5y=−30+−5y 15x=−5y−30 Step 2: Divide both sides by 15. 15x/15=−5y−30/15 x=−1/3 y−2

OpenStudy (mysticfury):

thanks @AloneS but i kinda need it in slope intercept form

alones (alones):

i think IT should hep then :3 https://www.desmos.com/

OpenStudy (mysticfury):

lol @AloneS i already tried that program XD

alones (alones):

Hrmm okay..i'm not sure then how should i help.. so you say you need this as intercept form?

OpenStudy (ijlal):

Do you know the standard form of a straight line equation is y=mx+c right?

OpenStudy (mysticfury):

yes but i need slope intercept form not standard form @ijlal

OpenStudy (ijlal):

just rearrange your equation in that form that is 15x+5y=-30 rearrange it into y=mx+c you will get ur intercept and slope :)

OpenStudy (mysticfury):

it dosent matter i didnt pass it anyway

OpenStudy (ijlal):

15x+5y=-30 5y=-15x-30 y=-6x-5 understand till here

OpenStudy (ijlal):

compare this equation with y=mx+c the term with x is the slope and the constant is the intercept

OpenStudy (ijlal):

@MysticFury

OpenStudy (ijlal):

and sorry a bit miscalculation by myself y=-6x-3 not y=-6x-5